# Controllability of the Fisher-Stefan system

### v1

@@ 4.2 Numerical experiments
This section provides numerical simulation results for both the nonlinear Fisher-Stephan problem (1.3) and its linearized approximation (system (2.1)) .
@@
4.2.1 Experiment 1: Linearized system (2.1)
This section shows numerical simulation results for system (2.1) and for the following data: $\Omega=(-1,1)$, $\overline{\psi}=0,\overline{h}=1$, ( then $m=0,n=-1$ ), $\mu=0.5$, $T=1$ and $f_{0}=g_{0}=0$. The initial conditions are $k_{0}=-0.5$ (so that $L_{0}=\sqrt{0.5})$, $z_{0}=\left\{\begin{array}[]{ll}\sin(\pi x),&0\leq x\leq 1,\\ 0,&-1\leq x<0,\end{array}\right.$ and the control region is $\omega=(-1+\varepsilon,-\varepsilon)$, with $\varepsilon=0.1$.
Table 1 collects results for the evolution, with respect to the number of training points, of all the losses in (4.1) as well as generalization error (difference between training and test errors). $20000$ epochs have been performed in all cases. Since our domain is rectangular, training points are the result of the Cartesian product of points in the spatial and time directions. The same number of random points has been taken in both directions. It is observed that beyond a certain number of training points (on the order of $10^{6}$), further increasing the number of points does not significantly enhance the numerical results.
Table 1: Experiment 1: Training errors for the different contributions in (4.1). Last column corresponds to generalization error ($\mathcal{E}_{\text{gener}}$).
@@
$N=10^{8}$ $6.0\times 10^{-4}$ $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ $2.0\times 10^{-4}$ $1.6\times 10^{-3}$ $1.0\times 10^{-5}$ $3.9\times 10^{-3}$ $4.5\times 10^{-2}$ $3.4\times 10^{-3}$
@@
Finally, Figure 4 (left) displays results for the controlled density $\varrho(xL(t),t)$ at several discrete times (left). Importantly, all the densities remain non-negative as a consequence of the penalization term $\mathcal{L}_{\hat{z}\geq 0}$ introduced in the loss function (4.1). The state variable $L(t)$ is plotted in Figure 4 (right). In both cases, it is observed that all the considered initial and controllability conditions are satisfied with accuracy.
4.2.2 Experiment 2: Nonlinear system (1.5)
This section reviews on numerical simulation results for system (1.5), which are obtained from the PINN approximation of (1.7). For comparison with its associated linearized system, we take the same initial and final conditions, namely, $h_{0}=0.5$, $\mu=0.5$, $T=1$, and $\psi_{0}(x)=\sin(\pi x)$.
@@
$N=10^{8}$ $7.4\times 10^{-3}$ $4.0\times 10^{-4}$ $3.0\times 10^{-4}$ $1.0\times 10^{-4}$ $1.1\times 10^{-3}$ $1.0\times 10^{-4}$ $1.8\times 10^{-3}$ $2.6\times 10^{-2}$ $7.5\times 10^{-3}$
@@
Figure 6: Experiment 2. $T=1$. (Left) Plot of the densities in problem (1.5) at different times, and (right) $L(t)$.
@@ 5 Conclusion
In this paper, we have analyzed the local exact controllability to the trajectories of the two components of the state variable in a one-dimensional free boundary problem governed by the Fisher-Stefan model. By applying a Neumann boundary control along the fixed boundary, we have shown that it is possible to steer the density towards that of an ”ideal” trajectory and to control the position of the free boundary. Furthermore, our results also establishes the null controllability of the system. The numerical approximation has been carried out using up-to-date, deep-learning-based approaches, specifically PINNs. A comparison study between the original nonlinear problem and its linearized version has been carried out. Although the pattern of the two computed controls are similar, quantitative differences are observed, which indicates that the nonlinear logistic term in the modeling of the problem really plays a role.
CRediT authorship contribution statement
Idriss Boutaayamou: Writing – review & editing, Supervision, Methodology, Investigation, Formal analysis. Fouad Et-tahri: Writing – review & editing, Writing – original draft, Investigation, Formal analysis, Visualization, Software. Lahcen Maniar: Writing – review & editing, Supervision, Methodology, Investigation, Formal analysis. Francisco Periago: Writing – review & editing, Methodology, Investigation, Formal analysis, Visualization, Software.
Declaration of competing interest
@@
Acknowledgements
F. Periago has been supported by grant PID2022-141957OA-C22 funded by
MCIN/AEI/10.13039/501100011033, by ”ERDF A way of making Europe”, and the Autonomous Community of the Región of Murcia, Spain, through the programme for the development of scientific and technical research by competitive groups (21996/PI/22), included in the Regional Program for the Promotion of Scientific and Technical Research of Fundación Séneca – Agencia de Ciencia y Tecnología de la Región de Murcia.
Data availability

### v2

@@ 4.2 Numerical experiments
This section provides numerical simulation results for both the nonlinear Fisher-Stephan problem (3) and its linearized approximation (system (14)) .
@@
4.2.1 Experiment 1: Linearized system (14)
This section shows numerical simulation results for Theorem 1 using linearized system (14) and for the following data: $\Omega=(-1,1)$, $\overline{\varrho}=0,\overline{L}=1$, ( then $\overline{\psi}=0,\overline{h}=1$, m=0,n=1m=0,n=-1 ), $\mu=0.5$, $T=1$ and $f_{0}=g_{0}=0$. The initial conditions are $k_{0}=-0.5$ (so that $L_{0}=\sqrt{0.5})$, $z_{0}=\left\{\begin{array}[]{ll}\sin(\pi x),&0\leq x\leq 1,\\ 0,&-1\leq x<0,\end{array}\right$. and the control region is $\omega=(-1+\varepsilon,-\varepsilon)$, with $\varepsilon=0.1$.
Table 1 collects results for the evolution, with respect to the number of training points, of all the losses in (38) as well as generalization error (difference between training and test errors). $20000$ epochs have been performed in all cases. Since our domain is rectangular, training points are the result of the Cartesian product of points in the spatial and time directions. The same number of random points has been taken in both directions. It is observed that beyond a certain number of training points (on the order of $10^{6}$), further increasing the number of points does not significantly enhance the numerical results.
Table 1: Experiment 1: Training errors for the different contributions in (38). Last column corresponds to generalization error ($\mathcal{E}_{\text{gener}}$).
@@ 4.2.1 Experiment 1: Linearized system (14)
$N=10^{8}$ $6.0\times 10^{-4}$ $\times 10^{-5} 1.0×10$$\times 10^{-5}$ $2.0\times 10^{-4}$ $1.6\times 10^{-3}$ $1.0\times 10^{-5}$ $3.9\times 10^{-3}$ $4.5\times 10^{-2}$ $3.4\times 10^{-3}$
@@
Finally, Figure 4 (left) displays results for the controlled density $\varrho(xL(t),t)$ at several discrete times (left). Importantly, all the densities remain non-negative as a consequence of the penalization term $\mathcal{L}_{\hat{z}\geq 0}$ introduced in the loss function (38). The state variable $L(t)$ is plotted in Figure 4 (right). In both cases, it is observed that all the considered initial and controllability conditions are satisfied with accuracy.
4.2.2 Experiment 2: Nonlinear system (5)
This section reviews on numerical simulation results for system (5), which are obtained from the PINN approximation of (7). For comparison with its associated linearized system, we take the same initial and final conditions, namely, $h_{0}=0.5$, $\mu=0.5$, $T=1$, and $\psi_{0}(x)=\sin(\pi x)$.
@@ 4.2.2 Experiment 2: Nonlinear system (5)
$N=10^{8}$ $7.4\times 10^{-3}$ $4.0\times 10^{-4}$ $3.0\times 10^{-4}$ $0\times 10^{-4} 1.1$$1\times 10^{-3}$ $1.0\times 10^{-4}$ $1.8\times 10^{-3}$ $2.6\times 10^{-2}$ $7.5\times 10^{-3}$
@@
Figure 6: Experiment 2. $T=1$. (Left) Plot of the densities in problem (5) at different times, and (right) $L(t)$.
@@ 5 Conclusion
In this paper, we have analyzed the local exact controllability to the trajectories of the two components of the state variable in a one-dimensional free boundary problem governed by the Fisher-Stefan model. By applying a Neumann boundary control along the fixed boundary, we have shown that it is possible to steer the density towards that of an "ideal" trajectory and to control the position of the free boundary. Furthermore, our results also establishes the null controllability of the system. The numerical approximation has been carried out using up-to-date, deep-learning-based approaches, specifically PINNs. A comparison study between the original nonlinear problem and its linearized version has been carried out. Although the pattern of the two computed controls are similar, quantitative differences are observed, which indicates that the nonlinear Logistic term in the modeling of the problem really plays a role.
[…]
Acknowledgements
F. Periago has been supported by grant PID2022-141957OA-C22 funded by MCIN/AEI/10.13039/501100011033, by "ERDF A way of making Europe", and the Autonomous Community of the Región of Murcia, Spain, through the programme for the development of scientific and technical research by competitive groups (21996/PI/22), included in the Regional Program for the Promotion of Scientific and Technical Research of Fundación Séneca – Agencia de Ciencia y Tecnología de la Región de Murcia.
We thank the anonymous referee for the valuable comments and suggestions.
Declarations